OpenStudy (anonymous):

A set of equations is given below: Equation C: y = 3x + 7 Equation D: y = 3x + 2 How many solutions are there to the given set of equations? A) One solution B) Two solutions C) Infinitely solutions D) No solution

OpenStudy (texaschic101):

one thing you should learn... in y = mx + b form, the slope is in the m position and the y intercept is in the b position. IF the slopes are the same and the y intercepts the same, then it is the same line and has infinite solutions. IF the slopes are the same and the y intercepts different, then it is a parallel line with no solutions. IF the slopes are different and the y intercepts different, then the equations will have 1 solution. So basically, what you have to do is find the slope and the y intercepts of both equations....compare them....then follow the steps that I put up above.

OpenStudy (texaschic101):

so do you know what the slope and the y intercept is for y = 3x + 7 ?? remember that in y = mx + b form, the slope is in the m position and the y intercept is in the b position

OpenStudy (anonymous):

no im really clue less

OpenStudy (texaschic101):

ok..let me walk you through this ...both of your equations are already in y = mx + b form....so we don't have to mess with that. In y = mx + b form, the slope is in the m position and the y intercept is in the b position. y = mx + b y = 3x + 7 so the number in the m position....3....is your slope and the number in the b position....7...is your y intercept are you understand this so far ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

now look at your other equation y = mx + b y = 3x + 2 slope(m) = 3 and y intercept (b) = 2 still following ?

OpenStudy (texaschic101):

now we are going to compare our slopes and y intercepts eq 1 : y = 3x + 7......slope(m) = 3 and y int (b) = 7 eq 2 : y = 3x + 2....slope(m) = 3 and y int (b) = 2 so we have equations with the same slope, but different y intercepts. And if you go back and read my first post, you will see that the lines are parallel, and therefore have 0 solutions